# Relative portfolio optimization via a value at risk based constraint

### v1

@@ 1. Introduction
Before we have a closer look at our results, let us mention the related literature. There have been a number of papers recently which consider strategic interactions between agents in asset allocation problems. The motivation stems from Brown et al., (2001) and Kempf and Ruenzi, (2008). The majority of papers model the interaction by maximizing the expected utility of relative wealth in financial markets of Black Scholes type. Relative wealth may be measured by a linear function of the terminal wealths of all agents, see e.g. Espinosa and Touzi, (2015) and Espinosa, (2010), or by a multiplicative function of the terminal wealths, see e.g. Basak and Makarov, (2014). For example Basak and Makarov, (2015) consider two agents investing in a Black Scholes model and maximizing the power utility of the ratio of both wealths. The difference between own wealth and arithmetic mean of other agents’ wealth has been treated in Espinosa and Touzi, (2015). Lacker and Zariphopoulou, (2019) investigate a model where each agent has her own financial market and consider the corresponding mean-field approach. This work has later been extended to investment-consumption in Lacker and Soret, (2020). Fu and Zhou, (2023) and Fu, (2023) provide a relation between Nash equilibria and systems of forward backward stochastic differential equations for agents with power utility and multiplicative relative performance for investment and consumption. Forward utilities with competition are for example considered in Musiela and Zariphopoulou, (2006), Dos Reis and Platonov, (2021) and Anthropelos et al., (2022). Models with more general financial market have been treated in Bäuerle and Göll, (2023), Kraft et al., (2020), Hu and Zariphopoulou, (2022), Aydoğan and Steffensen, (2024). Papers which model in addition price impacts of the investors are Bäuerle and Göll, (2024) and Curatola, (2024).
@@
On the other hand, there is a stream of literature considering benchmark and value-at-risk constraints for one agent problems. For example Basak and Shapiro, (2001) maximize expected utility under the constraint that, with a certain probability, the terminal wealth is above a given deterministic threshold. This is equivalent to requiring that the value-at-risk at the terminal time of the wealth is below a threshold. Gabih et al., (2005, 2009, 2006) consider a similar point of view replacing the value-at-risk by expected shortfall and Sass and Wunderlich, (2010) and Bäuerle and Chen, (2023) extend this situation to problems with partial information.
@@
In our model, we combine these two aspects. We consider $n$ agents who invest into the same financial market. We do not need a specific model for this market. It should be free of arbitrage and complete. The agents aim at maximizing their expected utility at terminal time $T$. Most of the paper is on logarithmic utility, but we also consider power utility. As a constraint, the probability that one’s wealth exceeds a linear combination of the other agents’ wealth is bounded from below. Thus, we adapt the model of Basak and Shapiro, (2001) to include relative concerns by replacing the constant solvency level in the optimization problem of agent $i$ by a weighted arithmetic mean of the other $n-1$ agents’ terminal wealth. Bell and Cover, (1980) consider a simple static two-person zero sum game where the payoff is the probability of beating the opponent’s outcome. There are also links to what is known as Colonel Blotto games, see Kovenock and Roberson, (2021). We incorporate the constraint
$\mathbb{P}\Big{(}X_{i}\geq\sum_{j\neq i}\beta_{j}X_{j}\Big{)}\geq\alpha_{i}$ (1.1)
@@
into the optimization problem of agent $i$, where $\alpha_{i}\in[0,1]$ and $\beta_{j}\in(0,1)$ for all $i,j\in\{1,\ldots,n\}$. By $X_{j}$ we denote the terminal wealth of agent $j\in\{1,\ldots,n\}$. Thus, in a fraction of $\alpha_{i}$ of the possible market scenarios, agent $i$ attains a terminal wealth which is at least as large as the weighted average of her competitors’ terminal wealth. The objective function of agent $i$ is given by the expected utility $\mathbb{E}U(X_{i})$ of her terminal wealth where $U$ is a CRRA utility. The parameters $\beta_{j}$ in (1.1) are custom to each agent $j$, but not to agent $i$, meaning that the weight assigned to agent $j$ is the same in the optimization of agent $i$ for any $i\neq j$. A possible choice is $\beta_{j}=\frac{1}{n-1}$, but we allow for more generality at this point. It is, for example, possible to consider weights in terms of the initial capital invested by the agents so that a larger initial investment goes along with a larger weight assigned to the corresponding agent. In contrast to the parameters $\beta_{j}$, $j=1,\ldots,n$, the level $\alpha_{i}\in[0,1]$ is chosen by agent $i$ herself. If $\alpha_{i}$ is chosen close to $1$, agent $i$ wants to insure her terminal wealth against the other agents’ wealth in almost all possible scenarios, while a value $\alpha_{i}$ close to 00 implies that she does not care as much about her performance with respect to the others. We look for Nash equilibria for this problem in terms of the final wealth. Depending on the specific parameters there are multiple or unique Nash equilibria. In the case of two agents, we obtain rather explicit results and are able to determine all Nash equilibria. In particular some discontinuity phenomena show up. In case of a logarithmic utility, as soon as some probabilities for the constraints are less than one, the structure of the classical optimal terminal wealth is enforced in the Nash equilibrium. The situation with many agents is considerably more difficult. We have to distinguish how large the probabilities for the constraints are. If the sum of all probabilities is at most one, the wealth constraints for the agents will be satisfied on disjoint events for all Nash equilibria.
@@
This paper is organized as follows. In the next section, we comment on the underlying financial market. Section 3 deals with the competition of two agents. We consider logarithmic and power utility there and derive all Nash equilibria rather explicitly. Section 4 is then dedicated to the $n$ agent case with $n\geq 3.$ In this section, we have to distinguish the cases $\alpha_{1}+\ldots+\alpha_{n}\leq 1$ and $\alpha_{1}+\ldots+\alpha_{n}>1.$ While we are able to determine the structure of all Nash equilibria in the first case, we restrict to searching for Nash equilibria of specific type in the second case. Section 5 deals with the special case of one agent aiming to beat a fixed stochastic benchmark with given probability. The last section discusses some numerical results and the appendix provides some auxiliary results.
@@ 2. Financial market
$Z_{T}=\exp\Big{(}-\gamma^{\top}W_{T}-\frac{1}{2}\|\gamma\|^{2}T\Big{)}$
with $\gamma=\sigma^{-1}\mu\in\mathbb{R}^{d}.$ Trading strategies are defined as $(\mathcal{F}_{t}^{W})-$ progressively measurable processes $(\pi_{t}^{k})$ denoting the amount of money invested at time $t$ in stock $k$. Since we compute the Nash equilibrium in terms of terminal wealths of the agents, we do not have to be specific about the financial market. The $\mathcal{F}$-measurable random variable $X_{i}$ always denotes the terminal wealth of agent $i$ at time $T$. Since the financial market is complete, all $\mathcal{F}$-measurable and integrable random variables can be attained by a certain portfolio strategy. The initial wealth needed for this portfolio is given by its price $\mathbb{E}[Z_{T}X_{i}].$ For more general semimartingale financial markets which fit into our setup see Bäuerle and Göll, (2023).
@@ 3.1. Logarithmic utility
We first consider the case where we have $n=2$ agents with a logarithmic utility. Both agents try to maximize their expected utility of terminal wealth $X_{i},i=1,2,$ at time $T$, given a fixed initial capital $x_{0}^{i}>0,i=1,2,$ under the constraint that their respective terminal wealth exceeds, with a certain probability, a fraction of the competitors terminal wealth. Since $T$ is fixed throughout, we delete it from the notation (except for $Z_{T}$). Thus, agent $i$ faces for fixed $X_{j},j\neq i$ with $i,j\in\{1,2\}$, the problem
@@
with $\alpha_{i},\beta_{i}\in[0,1]$ for $i=1,2$. The last equation ensures that the wealth $X_{i}$ has price $x_{0}^{i}$, i.e. can be attained by a self-financing strategy with initial wealth $x_{0}^{i}.$ Now we are seeking for a Nash equilibrium in this situation in terms of terminal wealths. Formally, a Nash equilibrium is here defined as follows.
@@
In what follows, we explicitly determine all Nash equilibria. In order to do this, we have to distinguish several parameter cases. Throughout we assume w.l.o.g. that $x_{0}^{1}\geq x_{0}^{2},$ i.e. agent 1 is at least as rich as agent 2.
@@
    If either $\alpha_{i}=1,\,i=1,2,\,\beta_{1}\beta_{2}<1$ or $\alpha_{1}<1,\,\alpha_{2}=1,$ the unique Nash equilibrium is given by
@@
    If $\alpha_{1}\leq 1,\,\alpha_{2}<1,$ there are two additional cases. If $x_{0}^{2}\geq\beta_{1}x_{0}^{1}$, the unique Nash equilibrium is given by
@@
    Note that in case a) of the previous theorem, the condition $x_{0}^{2}\geq\alpha_{2}\beta_{1}x_{0}^{1}$ already implies that $x_{0}^{1}=x_{0}^{2}$. Moreover, it also states that, as long as $x_{0}^{2}\geq\beta_{1}x_{0}^{1}$, the optimal solution $X_{i}^{*}=x_{0}^{i}Z_{T}^{-1}$, $i=1,2,$ to the unconstrained problem is always a Nash equilibrium. If at least one of the parameters $\alpha_{i},\beta_{i},\,i=1,2,$ is strictly smaller than $1$, it is the unique Nash equilibrium.

### v2

@@ 1. Introduction
Before we have a closer look at our results, let us mention the related literature. There have been a number of papers recently which consider strategic interactions between agents in asset allocation problems. The motivation stems from Brown et al., (2001) and Kempf and Ruenzi, (2008). The majority of papers model the interaction by maximizing the expected utility of relative wealth in financial markets of Black Scholes type. Relative wealth may be measured by a linear function of the terminal wealths of all agents, see e.g. Espinosa and Touzi, (2015) and Espinosa, (2010), or by a multiplicative function of the terminal wealths, see e.g. Basak and Makarov, (2014). More precisely, this means that $X_{T}^{i,\pi^{i}}-\frac{\theta_{i}}{n-1}\sum_{j\neq i}X_{T}^{j,\pi^{j}}$ or $X_{T}^{i,\pi^{i}}\big{(}\sum_{j\neq i}X_{T}^{j,\pi^{j}}\big{)}^{-\frac{\theta_{i}}{n-1}}$ enter the utility function of agent $i$ where $X_{T}^{i,\pi^{i}}$ is the wealth of agent $i$ at the terminal time $T$. For example, Basak and Makarov, (2015) consider two agents investing in a Black Scholes model and maximizing the power utility of the ratio of both wealths. The difference between own wealth and arithmetic mean of other agents’ wealth has been treated in Espinosa and Touzi, (2015). Lacker and Zariphopoulou, (2019) investigate a model where each agent has her own financial market and consider the corresponding mean-field approach. This work has later been extended to investment-consumption in Lacker and Soret, (2020). Fu and Zhou, (2023) and Fu, (2023) provide a relation between Nash equilibria and systems of forward backward stochastic differential equations for agents with power utility and multiplicative relative performance for investment and consumption. Forward utilities with competition are for example considered in Musiela and Zariphopoulou, (2006), Dos Reis and Platonov, (2021) and Anthropelos et al., (2022). Models with more general financial market have been treated in Bäuerle and Göll, (2023), Kraft et al., (2020), Hu and Zariphopoulou, (2022), Aydoğan and Steffensen, (2024). Papers which model in addition price impacts of the investors are Bäuerle and Göll, (2024) and Curatola, (2024).
@@
On the other hand, there is a stream of literature considering benchmark and value-at-risk constraints for one agent problems. For example Basak and Shapiro, (2001) maximize expected utility under the constraint that, with a certain probability, the terminal wealth is above a given deterministic threshold. This is equivalent to requiring that the value-at-risk at the terminal time of the wealth is below a threshold. Gabih et al., (2005, 2009, 2006) consider a similar point of view replacing the value-at-risk by expected shortfall and Sass and Wunderlich, (2010) and Bäuerle and Chen, (2023) extend this situation to problems with partial information. But in these models there is only a single agent and no competition involved.
@@
In our model, we combine these two aspects. We consider $n$ agents who invest into the same financial market. We do not need a specific model for this market. It should be free of arbitrage and complete. The agents aim at maximizing their expected utility at terminal time $T$. Most of the paper is on logarithmic utility, but we also consider power utility. As a constraint, the probability that one’s wealth exceeds a linear combination of the other agents’ wealth is bounded from below. Thus, we adapt the model of Basak and Shapiro, (2001) to include relative concerns by replacing the constant solvency level in the optimization problem of agent $i$ by a weighted arithmetic mean of the other $n-1$ agents’ terminal wealth. The motivation for such a model is threefold: First there is a psychological argument for private investors. Whether or not a return feels like a gain or loss is often relative to how other investors (friends, colleagues) performed or how the general stock market performed. This has for example been implemented in the cumulative prospect theory of Tversky and Kahneman, (1992) by introducing a reference point. Thus, investors try to beat such reference points. Second, fund managers often receive a part of their bonuses for outperforming other funds or benchmarks, see e.g. Browne, (1999). Thus, part of their attention is dedicated to outperform these benchmarks with at least a certain probability. Third, this can equivalently be considered as a portfolio optimization problem with a stochastic value-at-risk constraint, i.e. a value-at-risk constraint where the risk level is stochastic.
Bell and Cover, (1980) consider a simple static two-person zero sum game where the payoff is the probability of beating the opponent’s outcome. There are also links to what is known as Colonel Blotto games, see Kovenock and Roberson, (2021). We incorporate the constraint
$\mathbb{P}\Big{(}X_{i}\geq\sum_{j\neq i}\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}X_{j}\Big{)}\geq\alpha_{i}$ (1.1)
@@
into the optimization problem of agent $i$, where $\alpha_{i}\in[0,1]$ and $\beta_{ij}\in[0,1]$ for all $i,j\in\{1,\ldots,n\}$. By $X_{j}$ we denote the terminal wealth of agent $j\in\{1,\ldots,n\}$. Thus, in a fraction of $\alpha_{i}$ of the possible market scenarios, agent $i$ attains a terminal wealth which is at least as large as the weighted average of her competitors’ terminal wealth. The objective function of agent $i$ is given by the expected utility $\mathbb{E}U(X_{i})$ of her terminal wealth where $U$ is a CRRA utility. A possible choice is $\beta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}ij}}=\frac{1}{n-1}$, but we allow for more generality at this point. It is, for example, possible to consider weights in terms of the initial capital invested by the agents so that a larger initial investment goes along with a larger weight assigned to the corresponding agent. If $\alpha_{i}$ is chosen close to $1$, agent $i$ wants to insure her terminal wealth against the other agents’ wealth in almost all possible scenarios, while a value $\alpha_{i}$ close to 00 implies that she does not care as much about her performance with respect to the others. We look for Nash equilibria for this problem in terms of the final wealth. Depending on the specific parameters there are multiple or unique Nash equilibria. In the case of two agents, we obtain rather explicit results and are able to determine all Nash equilibria. In particular some discontinuity phenomena show up. In case of a logarithmic utility, as soon as some probabilities for the constraints are less than one, the structure of the classical optimal terminal wealth is enforced in the Nash equilibrium. The situation with many agents is considerably more difficult. We have to distinguish how large the probabilities for the constraints are. If the sum of all probabilities is at most one, the wealth constraints for the agents will be satisfied on disjoint events for all Nash equilibria.
@@
This paper is organized as follows. In the next section, we comment on the underlying financial market. Section 3 deals with the competition of two agents. We consider logarithmic and power utility there and derive all Nash equilibria rather explicitly. Section 4 is then dedicated to the $n$ agent case with $n\geq 3.$ In this section, we have to distinguish the cases $\alpha_{1}+\ldots+\alpha_{n}\leq 1$ and $\alpha_{1}+\ldots+\alpha_{n}>1.$ While we are able to determine the structure of all Nash equilibria in the first case, we restrict to searching for Nash equilibria of specific type in the second case. The last section discusses some numerical results and the appendix provides some auxiliary results.
@@ 2. Financial market
$Z_{T}=\exp\Big{(}-\theta^{\top}W_{T}-\frac{1}{2}\|\theta\|^{2}T\Big{)}$
with $\theta=\sigma^{-1}\mu\in\mathbb{R}^{d}.$ Trading strategies are defined as $(\mathcal{F}_{t}^{W})-$ progressively measurable processes $(\pi_{t}^{k})$ denoting the amount of money invested at time $t$ in stock $k$. Since we compute the Nash equilibrium in terms of terminal wealths of the agents, we do not have to be specific about the financial market. The $\mathcal{F}$-measurable random variable $X_{i}$ always denotes the terminal wealth of agent $i$ at time $T$. Since the financial market is complete, all $\mathcal{F}$-measurable and integrable random variables can be attained by a certain portfolio strategy. The initial wealth needed for this portfolio is given by its price $\mathbb{E}[Z_{T}X_{i}].$ For more general semimartingale financial markets which fit into our setup see Bäuerle and Göll, (2023).
Remark 2.1.
In order to motivate the discussion in the next sections, let us consider the following situation within the Black Scholes market with one stock, i.e. $d=1.$ We have only one agent who tries to invest in such a way that she maximizes her expected utility while beating (a fraction of) the stock price $S_{T}$ with at least a certain probability, i.e. our agent compares her wealth to the stock performance. More precisely, we consider the problem
$(P_{B})\left\{\begin{array}[]{l}\mathbb{E}\ln(X)\to\max\\ \mathbb{P}(X\geq\beta S_{T})\geq\alpha,\\ \mathbb{E}[Z_{T}X]=x_{0},\\ X\mbox{ is }\mathcal{F}-\mbox{measurable},\\ \end{array}\right.$
where $X$ is the terminal wealth of the agent. Thus, the agent maximizes the logarithmic utility of terminal wealth under the constraint that she beats at least with probability $\alpha$ the benchmark $\beta S_{T},\beta\in[0,1]$. The initial wealth is $x_{0}.$ Here we only state the optimal solution which can be derived with Lemma 7.2. We explain the derivation of results like this later, when we treat multi-agent models. Define
$\kappa(\lambda):=\Big{(}\frac{\lambda}{\beta}\Big{)}^{\frac{\mu}{\mu-\sigma^{2}}}\exp\Big{(}-\frac{1}{2}\mu T\Big{)},\quad\lambda\geq 0,$ (2.1)
and set $\lambda_{\alpha}$ such that $\mathbb{P}(\beta S_{T}\leq\lambda_{\alpha}/Z_{T})=\alpha.$ Then the optimal terminal wealth is $X^{*}=x_{0}/Z_{T}$ if $x_{0}\geq\lambda_{\alpha}$. Thus, if the agent is rich enough, the strategy which simply maximizes the expected utility will also satisfy the constraint. In the case $x_{0}<\lambda_{\alpha}$, we obtain
$X^{*}=\left\{\begin{array}[]{cl}\lambda/Z_{T},&\mbox{ if }Z_{T}\leq\kappa(\lambda_{\alpha})\mbox{ or }Z_{T}>\kappa(\lambda),\\ \beta S_{T},&\mbox{ if }Z_{T}\in(\kappa(\lambda_{\alpha}),\kappa(\lambda)],\end{array}\right.$
where $\lambda>\lambda_{\alpha}$ is such that $\mathbb{E}[X^{*}Z_{T}]=x_{0}.$ Thus, the terminal wealth is set equal to the reference $\beta S_{T}$ on a certain set of probability $\alpha$ and elsewhere the optimal wealth has the structure of the optimal wealth without constraint. This problem is related to a number of similar questions which have been treated in the literature before such as maximizing the probability of beating a benchmark or minimizing shortfall (see e.g. Browne,, 1999; Korn and Lindberg,, 2014; Föllmer and Schied,, 2016).
@@ 3.1. Logarithmic utility
We first consider the case where we have $n=2$ agents with a logarithmic utility. Both agents try to maximize their expected utility of terminal wealth $X_{i},i=1,2,$ at time $T$, given a fixed initial capital $x_{0}^{i}>0,i=1,2,$ under the constraint that their respective terminal wealth exceeds, with a certain probability, a fraction of the competitors terminal wealth. Since $T$ is fixed throughout, we delete it from the notation (except for $Z_{T}$). We also write $\beta_{1}$ instead of $\beta_{21}$ and $\beta_{2}$ instead of $\beta_{12}.$ Thus, agent $i$ faces for fixed $X_{j},j\neq i$ with $i,j\in\{1,2\}$, the problem
@@
with $\alpha_{i},\beta_{i}\in[0,1]$ for $i=1,2$. The last equation ensures that the wealth $X_{i}$ has price $x_{0}^{i}$, i.e. can be attained by a self-financing strategy with initial wealth $x_{0}^{i}.$ Throughout we assume that all parameters, preferences and the opportunity investment set are known to all agents. Now we are seeking for a Nash equilibrium in this situation in terms of terminal wealths. Formally, a Nash equilibrium is here defined as follows.
@@
In what follows, we explicitly determine all Nash equilibria. Throughout we assume w.l.o.g. that $x_{0}^{1}\geq x_{0}^{2},$ i.e. agent 1 is at least as rich as agent 2. To obtain the Nash equilibria we have to solve the best response problems $(P_{2})$ for both agents and then find a fixed point. The best response problems are utility maximization problems with side constraints of the form $X_{i}\geq Y\mathbbm{1}_{A}$ with $\mathbb{P}(A)=\alpha_{i}$ where the location of the set $A$ varies. Given that $\lambda/Z_{T}$ is the form of the optimal wealth in the log-utility problem without constraint, it is not too hard to guess that the optimal wealth with constraint satisfies $X_{i}=\max\{Y,\lambda/Z_{T}\}$ on a set of probability $\alpha$ and elsewhere takes the form $\lambda/Z_{T}$. The crucial question is where the set $A$ is located. We discuss the solution of the best response problem in detail in the appendix. In Lemma 7.2 in Appendix 7.2 we show that the set $A$ is located where it is cheapest to deviate from $\lambda/Z_{T}$ to satisfy the constraint, i.e. we increase $\lambda$ such that the set $\{\lambda/Z_{T}\geq Y\}$ has probability $\alpha.$ We now have to distinguish several parameter cases.
@@
    If $\alpha_{2}=1$ and $\alpha_{1}\beta_{1}\beta_{2}<1$ then the unique Nash equilibrium is given by
@@
    If $\alpha_{1}\leq 1,\,\alpha_{2}<1,$ there are two cases. If $x_{0}^{2}\geq\beta_{1}x_{0}^{1}$, the unique Nash equilibrium is given by
@@
    Note that in case a) of the previous theorem, the condition $x_{0}^{2}\geq\alpha_{2}\beta_{1}x_{0}^{1}$ already implies that $x_{0}^{1}=x_{0}^{2}$. In parts b) and c) we see that as long as $x_{0}^{2}\geq\beta_{1}x_{0}^{1}$, the optimal solution $X_{i}^{*}=x_{0}^{i}Z_{T}^{-1}$, $i=1,2,$ to the unconstrained problem is always a Nash equilibrium, independent of $\alpha_{i}$. This is because comparing the $X_{i}^{*}$ boils down to comparing the constants and the constraint is obviously satisfied with probability 1. This also makes sense from an economic point of view, because this is the case where the constraint is already satisfied for the initial wealths and due to the ’no arbitrage’ condition it carries over to the terminal wealths. As soon as $x_{0}^{2}<\beta_{1}x_{0}^{1}$ we always get different Nash equilibria since the constraint is satisfied with probability 0 for the unconstraint solutions.
@@
3.2. Replicating strategies for the Nash equilibrium
Until now, we have only discussed the optimal terminal wealth in the Nash equilibrium. However, since we assumed that the financial market is complete, there exist replicating portfolio strategies for the Nash equilibrium from Theorem 3.3. We use the Black-Scholes market discussed in Section 2 in order to give explicit representations of the replicating strategies. In case b) of Theorem 3.3, the replicating strategies in terms of fractions of wealth invested into the stock are known to be constant (Mertion ratio, see e.g. Merton,, 1969; Korn,, 1997) and given by $(\sigma\sigma^{\top})^{-1}\mu$. For later comparison, we consider the process $\pi_{2}^{B}$ describing the invested amount of wealth (instead of the fraction), which is given by
$\pi_{2}^{B}(t)=\frac{x_{0}^{2}}{Z_{t}}\big{(}\sigma\sigma^{\top}\big{)}^{-1}\mu.$
Thus, we only discuss the replicating strategy of the second investor in case c) of Theorem 3.3. Moreover, we cannot give an explicit representation for arbitrary sets $A_{2}$ and therefore only consider sets of the form $A_{2}=\{c_{1}<Z_{T}<c_{2}\}$ for constants $0\leq c_{1}<c_{2}\leq\infty$ with $\mathbb{P}(c_{1}<Z_{T}<c_{2})=\alpha_{2}$. Then we can use Lemma 7.3 from Appendix 7.4 to find the portfolio-wealth pair $(X_{2}^{*}(t),\pi_{2}^{*}(t))_{t\in[0,T]}$ replicating $X_{2}^{*}$. To simplify notation, we introduce the abbreviation
[…]
Here, $(Z_{t})_{t\in[0,T]}$ denotes the state price density process, i.e.
$Z_{t}=\exp\Big{(}-\theta^{\top}W_{t}-\frac{1}{2}\|\theta\|^{2}t\Big{)},\,t\in[0,T].$
Theorem 3.5.
The replicating portfolio-wealth pair for $X_{2}^{*}$ from Theorem 3.3 c) is given by
$X_{2}^{*}(t)$ $=\frac{x_{0}^{2}-\alpha_{2}\beta_{1}x_{0}^{1}}{1-\alpha_{2}}\Big{(}Y_{1}(t)+Y_{3}(t)\Big{)}+\beta_{1}x_{0}^{1}Y_{2}(t),$
$\pi_{2}^{*}(t)$ $=\frac{x_{0}^{2}-\alpha_{2}\beta_{1}x_{0}^{1}}{1-\alpha_{2}}\Big{(}\varphi_{1}(t)+\varphi_{3}(t)\Big{)}+\beta_{1}x_{0}^{1}\varphi_{2}(t),$
where
$Y_{1}(t)$ $=\frac{1}{Z_{t}}\Phi\big{(}f(c_{1},t)\big{)},\quad Y_{2}(t)=\frac{1}{Z_{t}}\big{[}\Phi\big{(}f(c_{2},t)\big{)}-\Phi\big{(}f(c_{1},t)\big{)}\big{]},\quad Y_{3}(t)=\frac{1}{Z_{t}}\Phi\big{(}-f(c_{2},t)\big{)},$
$\varphi_{1}(t)$ $=\Bigg{[}Y_{1}(t)+\frac{1}{Z_{t}\|\theta\|\sqrt{T-t}}\varphi\big{(}f(c_{1},t)\big{)}\Bigg{]}\big{(}\sigma\sigma^{\top}\big{)}^{-1}\mu,$
$\varphi_{2}(t)$ $=\Bigg{[}Y_{2}(t)+\frac{1}{Z_{t}\|\theta\|\sqrt{T-t}}\big{(}\varphi\big{(}f(c_{2},t)\big{)}-\varphi\big{(}f(c_{1},t)\big{)}\big{)}\Bigg{]}\big{(}\sigma\sigma^{\top}\big{)}^{-1}\mu,$
$\varphi_{3}(t)$ $=\Bigg{[}Y_{3}(t)-\frac{1}{Z_{t}\|\theta\|\sqrt{T-t}}\varphi\big{(}f(c_{2},t)\big{)}\Bigg{]}\big{(}\sigma\sigma^{\top})^{-1}\mu.$
Here the portfolio strategy $\pi_{2}^{*}$ denotes the amount of wealth invested into the stocks at time $t\in[0,T).$
Proof.
If we write $X_{2}^{*}$ from Theorem 3.3 c) for $A_{2}=\{c_{1}<Z_{T}<c_{2}\}$ as
$X_{2}^{*}=\frac{x_{0}^{2}-\alpha_{2}\beta_{1}x_{0}^{1}}{1-\alpha_{2}}Z_{T}^{-1}\left(\mathbbm{1}\{Z_{T}\leq c_{1}\}+\mathbbm{1}\{Z_{T}\geq c_{2}\}\right)+\beta_{1}x_{0}^{1}Z_{T}^{-1}\mathbbm{1}\{c_{1}<Z_{T}<c_{2}\},$
the result follows from Lemma 7.3 and the linearity of the wealth process. ∎